Will someone help me solve this problem or check to see if I am on the right track? I think I would start by dividing 15/2 to get a unit rate of 7.50 per book and use $90 as a starting rate. M=90-7.50b
So far 0 in the linear table, it would be M=90-7.50*0=90
Noah had 90 dollars spending money from a math competition he won. He found some math problem books that cost $15.00 for every 2 books. Of course, he decided to buy many of them! a. Write a linear equation to model this situation using M = Total Money River has left and B = Number of
books bought and complete the table.

Ind. Variable Dep. Variable
B M
0
1
2

Answer by math_tutor2020(3817)


You have the correct equation M = 90-7.50b which is the same as M = -7.50b+90

From here you plug in values b = 0, b = 1, b = 2 to get corresponding values of M

You've already shown that b = 0 leads to M = 90.

You repeat for b = 1
M = 90-7.5b
M = 90-7.5*1
M = 82.50
Telling us that buying 1 book means Noah has $82.50 left

And for b = 2 also
M = 90-7.5b
M = 90-7.5*2
M = 75
Showing that buying 2 books means he has $75 left

The table looks like this
BM
090
182.50
275


So it seems like you're on the right track.